# Balancing Embedding Spectrum for Recommendation

### v1

@@ V-A1 Datasets
  • CiteULike [^2] https://github.com/js05212/citeulike-a: This is a scholarly article recommendation dataset. Users are allowed to create their own collections of articles including abstracts, titles, and tags, etc.
  • Yelp [27]: This is a business dataset from Yelp Challenge data. The items are point of interests (POIs), users can leave reviews and ratings.
  • Gowalla [15]: This is a check-in datasets recording which locations users have visited.
@@
V-A2 Evaluation Metrics
We adopt two widely used evaluation metrics for personalized rankings: Recall and normalized discounted cumulative gain (nDCG) [28] to evaluate model performance. The recommendation list is generated by ranking unobserved items and truncating at position $k$. Recall measures the ratio of the relevant items in the recommendation list to all relevant items in test sets, while nDCG takes the rank into consideration by assigning higher scores to the relevant items ranked higher. We use 80% of the interactions for training and randomly select 10% from the training set as validation set for hyper-parameter tuning, the rest 20% data is used for test sets; we report the average accuracy on test sets.
V-A3 Baselines
@@
  • BPR [12]: This is a stable and classic MF-based method, exploiting a Bayesian personalized ranking loss for personalized rankings.
  • CL-Rec [22]: This is a CL-based method. Since there are no item features on the datasets used in this work, we remove data augmentation and use BPR as the main loss to better compare with our methods.
@@
  • ENMF [29]: This is a neural recommendation model performing whole-data-based learning without sampling. We choose user-based ENMF as the baseline.
@@
  • CCL [13]: The proposed cosine contrastive loss (CCL) maximizes the similarity between positive pairs and minimizes the similarity of negative pairs below the margin $m$. After parameter tuning, we set $\mathcal{N}=1000$, $w=300$ on all datasets, $m=0.1$, $0.3$, and $0.3$ on CiteULike, Yelp, and Gowalla, respectively, and we choose MF as the encoder.
@@
  • DirectAU [23]: This method directly optimizes alignment and uniformity and shows superior performance. Following the original paper, we choose MF and LightGCN as the encoder, and set $\gamma=5.0$, 0.5, and 1.5 on CiteULike, Yelp, and Gowalla, respectively.
@@
  • LightGCN [14]: This is a linear GNN method that only keeps neighborhood aggregation for recommendation. We employ a three-layer architecture as our baseline.
@@
  • SGL-ED [21]: This model explores self-supervised learning by maximizing the agreements of multiple views from the same node, where the node views are generated by adding noise such as randomly removing the edges or nodes on the original graph. We set $\tau=0.2$, $\lambda_{1}=0.1$, $p=0.1$, and use a three-layer architecture.
@@
  • LightGCL [30]: This is a simple yet effective graph contrastive learning method injecting global collaborative relations via singular value decomposition.
@@
  • GDE [31]: This method only keeps a very few graph features for recommendation without stacking layers, showing less complexity and higher efficiency than conventional GCN-based methods.
@@
  • LogDet [32]: This is a decorrelation-enhanced method to mitigate the dimensional collapse of graph collaborative filtering.
Figure 5: The extent of collapse on three datasets.
Figure 6: All models are optimized without using observed interactions on Yelp. We show how erank changes on DirectSpec (MF) and MF (BCE) where $N$ is the negative sampling ratio in (a). We use BCE (MF) for the first 50 epochs and DirectSpec for the last 50 epochs in (b).
V-A4 Implementation Details
@@
We implemented our DirectSpec based on PyTorch and the code is released publicly [^3] https://github.com/tanatosuu/directspec. SGD is adopted as the optimizer for all models, the embedding size $d$ is set to 64, the regularization rate is set to 0.01 on all datasets, the learning rate is tuned amongst $\{0.001,0.005,0.01,\cdots\}$; without specification, the model parameters are initialized with Xavier Initialization [33]; the batch size is set to 256. We report other hyperparameter settings in the next subsection.
TABLE V: Training time (seconds) per epoch.
@@
V-B Comparison
V-B1 Performance
We report overall performance in Table IV, and observe the followings:
  • Overall, GCN-based methods show better performance on sparse data (Yelp and Gowalla) than dense data (CiteULike). For instance, DirectSpec+(GCN) achieves better and worse than DirectSpec+(MF) on Gowalla and CiteULike, respectively, and their performance is close on Yelp. LightGCN underperforms BPR on CiteULike and Gowalla, which might be attributed to the slow convergence that has been reported in the original paper. Among baselines, DirectAU achieves the best on CiteULike and Yelp, while LightGCL outperforms other baselines on Gowalla. Our proposed DirectSpec+ implemented on both LightGCN and MF consistently show improvements over all baselines.
  • CL-Rec is even inferior to BPR, indicating that it is hard for traditional CF task without any side information to benefit from CL. In the meanwhile, the GCL-based method (SGL-ED and LightGCL) works much better as we can exploit the graph structure containing rich topological information as the input, as opposed to traditional CF task that only uses user/item ID as input.
  • DirectAU directly regulating the uniformity shows relatively superior performance. Inspired by [20], it adopts a Radial Basis Function kernel which shares a similar form to the uniformity loss. Since we have shown that the uniformity loss can tackle embedding collapse by indirectly balancing the spectrum distribution in Section IV-D1, the effectiveness of DirectAU demonstrates our previous analysis.
  • Since our methods are implemented on MF and LightGCN, the effectiveness of our methods can be further demonstrated by directly comparing with them. DirectSpec(MF) and DirectSpec+(MF) outperform BPR by 14.3% and 35.8%, on average in terms of nDCG@10, respectively. In the meanwhile, the improvement of DirectSpec(GCN) and DirectSpec+(GCN) over LightGCN is 15.7% and 32.9% on average in terms of nDCG@10, respectively.
@@ V-B1 Performance
  • DirectSpec+ shows significant improvement over DirectSpec across all datasets. As introduced in Section IV-C DirectSpec+ is more effective tackling embedding collapse than DirectSpec by effectively penalizing the hard negative samples with user/item pairs highly correlated. The superior performance demonstrates its effectiveness.
@@
(b) Gowalla
(c) Yelp
Figure 7: Erank comparison among several methods.
@@
(c) Erank comparison
Figure 8: How DirectSpec prevents collapse (on Yelp).
V-B2 Efficiency
We report the training time per epoch of BPR, LightGCN, and DirectSpec+ (MF and GCN) in Table V. The experiments are all conducted on Intel(R) Core(TM) i9-10980XE CPU and NVIDIA RTX A6000 GPU. We first compare BPR and DirectSpec+(MF), and observe that it takes roughly 0.5x additional running time, which is acceptable considering the significant improvement. DirectSpec+(GCN) almost shows no additional running time over LightGCN, due to the reason that the complexity of DirectSpec+ is mainly determined by the batch size rather than the size of users and items. Since LightGCN has a higher complexity than BPR, the proportion of DirectSpec+(LightGCN)’s running time to that of LightGCN is much smaller than the proportion of DirectSpec+(MF)’s running time to that of BPR. This finding can explain another observation that the additional time is smaller on larger datasets (i.e., Gowalla$<$Yelp$<$CiteULike). The complexity of DirectSpec+ is unchanged while LightGCN and BPR accordingly take more running time on larger datasets, thus the proportion of DirectSpec+’s running time to that of LightGCN/BPR becomes smaller. Meanwhile, SGL-ED is more computationally expensive on larger datasets, and DirectSpec+(LightGCN) is more efficient than it with more significant improvement over LightGCN.
V-B3 Comparison w.r.t Effective Rank
@@ V-B2 Efficiency
Negative Sampling. The results in Table IV demonstrates the superiority of DirectSpec over negative sampling based methods including BPR and CCL in terms of performance. Furthermore, we compare how these two methods alleviate collapse issue by optimizing them solely based on unobserved interactions. In Figure 6 (a), we observe that the erank of MF (BCE) consistently drops without optimization on the observed interactions, resulting in a complete collapse, and the erank drops more quickly as increasing the negative sampling ratios, whereas the erank of DirectSpec remains unchanged throughout the training. This observation indicates that there exists a dynamic balance between optimization on observed and unobserved interactions, the lack of either of them cannot prevent the embeddings from collapsing. Furthermore, since negative sampling is equivalent to a high pass filter, it can only alleviate the collapse caused by low pass filters (e.g., optimization on observed interactions), while DirectSpec is equivalent to an all pass filter that can tackle the collapse under any situations. Figure 6 (b) demonstrates that DirectSpec can recover the embeddings from collapsing caused by negative sampling.
LightGCL is an effective and efficient CL-based method with the design of uniformity helping alleviate the collapse issue. As shown in Figure 7 (a), the erank of LightGCL is lower than DirectSpec+ at the beginning, gradually increases as training proceeds and converges to a value smaller than the erank of DirectSpec+, demonstrating the superiority of DirectSpec tackling the collapse issue over LightGCL, explaining why it shows better performance.
LogDet is an enhanced method improving based on uniformity-based methods. As shown in Figure 7 (b), the erank of LogDet remains stable and closer to $d$ than LightGCL, showing a better ability tackling collapse issue than uniformity-based methods. Meanwhile, it still exhibits a lower erank than our proposed DirectSpec+ throughout the training.
Solutions to Over-Smoothing. As shown in Section IV, embedding collapse is attributed to the optimization part and over-smoothing (for GNN-based methods). In this subsection, we investigate (1) if the solutions to over-smoothing can also well address the collapse issue and (2) Do DirectSpec show superiority over them in addressing the over-smoothing. To this end, we compare DropEdge [25] (implemented on LightGCN), a solution to over-smoothing, LightGCN (suffer from both (1) and (2)), and BPR (suffer from (1)) with DirectSpec+. We report how their eranks change throughout the training in Figure 7 (c). We can see the erank of DropEdge is higher than LightGCN while lower than BPR during the training, indicating that the solution to over-smoothing can only alleviate (2) while fails to tackle (1), the main cause of collapse issue analyzed in this work. Furthermore, DirectSpec+ shows a higher and more stable erank than DropEdge, demonstrating a better capability in addressing over-smoothing than DropEdge.
V-C Model Analysis
V-C1 Can DirectSpec Prevent Collapse?
In Figure 8 (a), we report how the erank changes as training proceeds on BPR, LightGCN, and DirectSpec+(MF). For BPR and LightGCN, the erank plunges significantly in the first place, then increases slowly and tends to converge to a value smaller than $d$, resulting in an incomplete embedding collapse. While the erank of DirectSpec+(MF) is close to $d$ at the beginning of training and remains almost unchanged throughout the training. Figure 8 (b) illustrates the normalized singular value distribution of the embedding matrix when the accuracy of BPR and DirectSpec+(MF) is maximized. We can see that most singular values of BPR’s embeddings are less than 0.2, barely contributing to the representations; while DirectSpec+(MF)’s singular values are mostly larger than 0.5, indicating that the representations are contributed by more dimensions. One may raise a concern, that DirectSpec still suffers from embedding collapse as the erank is still slightly smaller than $d$ ($d=64>erank(\mathbf{H})>62$ in our experiments) when the accuracy is maximized. As shown in Section IV-B, the goal of DirectSpec is equivalent to decorrelating users and items, while it is obvious that not all users/items are irrelevant, considering the relations of observed interactions (e.g, users share similar/opposite interests are positively/negatively related; items that are interacted by similar users or users that have opposite interests are also positively or negatively correlated, respectively). Thus, a higher erank does not necessarily lead to a better accuracy. From a spectral perspective, all dimensions are not completely equally important to users/item. Some dimensions might contain more important information while others contain less. In other words, the goal of DirectSpec is to assure that all dimensions literally contribute to the representations, as opposed to the representations of BPR or LightGCN that are only predominantly distributed along a few dimensions. In Figure 8 (c), DirectSpec seems to show a better ability of balancing spectrum as it has a higher erank, while it shows inferior performance to DirectSpec+. As shown in Equation (15), all samples are decorrelated with the same pace on DirectSpec, thus the similar or dissimilar (i.e., sharing opposite interests) users/items that should be correlated to some extent are also perfectly orthogonalized, leading to underfitting.

### v2

@@ 5.1.1. Datasets
  • CiteULike [^1] https://github.com/js05212/citeulike-a: This is an scholarly article recommendation dataset. Users are allowed to create their own collections of articles including abstracts, titles, and tags, etc.
  • Yelp (He et al., 2016): This is a business dataset from Yelp Challenge data. The items are point of interests (POIs), users can leave reviews and ratings.
  • Gowalla (Wang et al., 2019): This is a check-in datasets recording which locations users have visited.
@@
5.1.2. Evaluation Metrics
We adopt two widely used evaluation metrics for personalized rankings: Recall and normalized discounted cumulative gain (nDCG) (Järvelin and Kekäläinen, 2002) to evaluate model performance. The recommendation list is generated by ranking unobserved items and truncating at position $k$. Recall measures the ratio of the relevant items in the recommendation list to all relevant items in test sets, while nDCG takes the rank into consideration by assigning higher scores to the relevant items ranked higher. We use 80% of the interactions for training and randomly select 10% from the training set as validation set for hyper-parameter tuning, the rest 20% data is used for test sets; we report the average accuracy on test sets.
5.1.3. Baselines
@@
  • BPR (Rendle et al., 2009): This is a stable and classic MF-based method, exploiting a Bayesian personalized ranking loss for personalized rankings.
@@
  • ENMF (Chen et al., 2020b): This is a neural recommendation model performing whole-data-based learning without sampling. We choose user-based ENMF as the baseline.
@@
  • CCL (Mao et al., 2021): The proposed cosine contrastive loss (CCL) maximizes the similarity between positive pairs and minimizes the similarity of negative pairs below the margin $m$. After parameter tuning, we set $\mathcal{N}=1000$, $w=300$ on all datasets, $m=0.1$, $0.3$, and $0.3$ on CiteULike, Yelp, and Gowalla, respectively, and we choose MF as the encoder.
@@
  • DirectAU (Wang et al., 2022): This method directly optimizes alignment and uniformity and shows superior performance. Following the original paper, we choose MF and LightGCN as the encoder, and set $\gamma=5.0$, 0.5, and 1.5 on CiteULike, Yelp, and Gowalla, respectively.
@@
  • LightGCN (He et al., 2020): This is a linear GNN method that only keeps neighborhood aggregation for recommendation. We employ a three-layer architecture as our baseline.
@@
  • SGL-ED (Wu et al., 2021): This model explores self-supervised learning by maximizing the agreements of multiple views from the same node, where the node views are generated by adding noise such as randomly removing the edges or nodes on the original graph. We set $\tau=0.2$, $\lambda_{1}=0.1$, $p=0.1$, and use a three-layer architecture.
@@
  • LightGCL (Cai et al., 2023): This is a simple yet effective graph contrastive learning method injecting global collaborative relations via singular value decomposition.
@@
  • GDE (Peng et al., 2022a): This method only keeps a very few graph features for recommendation without stacking layers, showing less complexity and higher efficiency than GNN-based methods.
@@
  • LogDet (Zhang et al., 2023): This is a decorrelation-enhanced method to mitigate the dimensional collapse of graph collaborative filtering.
  • AFDGCF (Wu et al., 2024): This method tackles the over-correlation issue in GNN-based methods, and shares significant similarities with Barlow Twin (Zbontar et al., 2021). It can be considered as the implementation of Barlow Twin for recommendation. We choose LightGCN as the encoder.
5.1.4. Implementation Details
@@
We implemented our DirectSpec based on PyTorch, and the code is released publicly [^2] https://github.com/tanatosuu/directspec. SGD is adopted as the optimizer for all models, the embedding size $d$ is set to 64, the regularization rate is set to 0.01 on all datasets, the learning rate is tuned amongst $\{0.001,0.005,0.01,\cdots\}$; without specification, the model parameters are initialized with Xavier Initialization (Glorot and Bengio, 2010); the batch size is set to 256. We report other hyperparameter settings in the next subsection.
Table 3. Performance comparison. The best baseline is underlined. “*” indicates statistical significance at $p<0.01$ for a one-tailed t-test.
Datasets Methods nDCG@10 nDCG@20 Recall@10 Recall@20
CiteULike BPR 0.1620 0.1773 0.1778 0.2190
CCL(MF) 0.1545 0.1642 0.1716 0.1996
ENMF-U 0.1335 0.1601 0.1556 0.2193
DirectAU(MF) 0.2102 0.2252 0.2260 0.2693
DirectAU(GCN) 0.1926 0.2109 0.2116 0.2604
LightGCN 0.1610 0.1781 0.1771 0.2190
SGL-ED 0.1890 0.2065 0.2117 0.2588
LightGCL 0.2096 0.2238 0.2214 0.2638
GDE 0.1890 0.2061 0.2055 0.2528
LogDet 0.1403 0.1579 0.1548 0.2034
AFDGCF 0.1619 0.1787 0.1790 0.2246
DirectSpec(MF) 0.1688 0.1849 0.1827 0.2270
DirectSpec(GCN) 0.1693 0.1863 0.1875 0.2334
DirectSpec+(MF) 0.2197 0.2354 0.2352 0.2818
DirectSpec+(GCN) 0.2038 0.2213 0.2213 0.2704
Yelp BPR 0.0487 0.0583 0.0607 0.0869
CCF(MF) 0.0509 0.0593 0.0617 0.0846
ENMF-U 0.0665 0.0811 0.0786 0.1167
DirectAU(MF) 0.0721 0.0848 0.0872 0.1219
DirectAU(GCN) 0.0695 0.0819 0.0854 0.1194
LightGCN 0.0572 0.0690 0.0721 0.1045
SGL-ED 0.0676 0.0794 0.0837 0.1166
LightGCL 0.0673 0.0790 0.0836 0.1151
GDE 0.0653 0.0771 0.0805 0.1129
LogDet 0.0711 0.0833 0.0863 0.1203
AFDGCF 0.0585 0.0702 0.0743 0.1065
DirectSpec(MF) 0.0689 0.0804 0.0839 0.1156
DirectSpec(GCN) 0.0712 0.0832 0.0861 0.1192
DirectSpec+(MF) 0.0743 0.0864 0.0909 0.1249
DirectSpec+(GCN) 0.0745 0.0868 0.0909 0.1252
Gowalla BPR 0.1164 0.1255 0.1186 0.1483
CCL(MF) 0.1269 0.1349 0.1295 0.1573
ENMF-U 0.1166 0.1286 0.1173 0.1528
DirectAU(MF) 0.1286 0.1402 0.1349 0.1710
DirectAU(GCN) 0.1298 0.1409 0.1363 0.1711
LightGCN 0.0987 0.1098 0.1074 0.1399
SGL-ED 0.1343 0.1462 0.1417 0.1779
LightGCL 0.1368 0.1482 0.1436 0.1801
GDE 0.1261 0.1367 0.1313 0.1656
LogDet 0.1179 0.1272 0.1203 0.1511
AFDGCF 0.1011 0.1117 0.1087 0.1406
DirectSpec(MF) 0.1133 0.1225 0.1183 0.1480
DirectSpec(GCN) 0.1160 0.1251 0.1191 0.1490
DirectSpec+(MF) 0.1389 0.1509 0.1447 0.1819
DirectSpec+(GCN) 0.1405 0.1518 0.1467 0.1837
Table 4. Training time (seconds) per epoch.
@@
AFDGCF(GCN) 11.45 (1.30x) 48.53 (1.12x) 84.86 (1.06x)
5.2. Comparison
5.2.1. Performance
We report overall performance in Table 3, and observe the followings:
  • Overall, GNN-based methods show better performance on sparse data (Yelp and Gowalla) than dense data (CiteULike). For instance, DirectSpec+(GCN) achieves better and worse than DirectSpec+(MF) on Gowalla and CiteULike, respectively, and their performance is close on Yelp. LightGCN underperforms BPR on CiteULike and Gowalla, which might be attributed to the slow convergence that has been reported in the original paper. Among baselines, DirectAU achieves the best on CiteULike and Yelp, while LightGCL outperforms other baselines on Gowalla. Our proposed DirectSpec+ implemented on both LightGCN and MF consistently show improvements over all baselines.
  • ENMF and CCL exploit enhanced negative sampling for training, while the superior performance over them demonstrates the superiority of DirectSpec learning paradigm.
  • Unlike negative sampling, SSL-based methods adopt different training strategies to enhance the data representations. Amongst them, DirectAU directly regulating the uniformity shows relatively superior performance, while AGDGCF utilizing a Barlow-Twin-like model to remove the feature correlation shows poor performance.
  • Since our methods are implemented on MF and LightGCN, the effectiveness of our methods can be further demonstrated by directly comparing with them. DirectSpec(MF) and DirectSpec+(MF) outperform BPR by 14.3% and 35.8%, on average in terms of nDCG@10, respectively. In the meanwhile, the improvement of DirectSpec(GCN) and DirectSpec+(GCN) over LightGCN is 15.7% and 33.1% on average in terms of nDCG@10, respectively.
@@ 5.2.1. Performance
  • DirectSpec+ shows significant improvement over DirectSpec across all datasets. As introduced in Section 4.3, we adopt a individualized spectrum balancing policy for DirectSpec+ to achieve a better trade-off between alignment and spectrum balancing. The superior performance demonstrates its effectiveness.
Figure 5. The extent of collapse on three datasets.
@@
(b) Yelp
(c) CiteULike
(d) Yelp
Figure 6. All models are optimized without using observed interactions. In (a) and (b), we show how erank changes on DirectSpec (MF) and MF (BCE) where $N$ is the negative sampling ratio. In (c) and (d), we use BCE (MF) for the first 50 epochs and DirectSpec for the last 50 epochs.
5.2.2. Efficiency
We report the training time per epoch of several methods implemented on BPR and LightGCN in Table 4. The experiments are all conducted on Intel(R) Core(TM) i9-10980XE CPU and NVIDIA RTX A6000 GPU. We first compare BPR and DirectSpec+(MF), and observe that it takes roughly 0.5x additional running time, which is acceptable considering the significant improvement. Compared to other methods implemented on LightGCN, DirectSpec+ almost shows no additional running time, due to the reason that the complexity of DirectSpec+ is mainly determined by the batch size rather than the size of users and items. Since LightGCN has a higher complexity than BPR, the proportion of DirectSpec+(LightGCN)’s running time per epoch to that of LightGCN is much smaller than the proportion of DirectSpec+(MF)’s running time to that of BPR. This finding can explain another observation that the additional time is smaller on larger datasets (i.e., Gowalla$<$Yelp$<$CiteULike). The complexity of DirectSpec+ is unchanged while LightGCN and BPR accordingly take more running time on larger datasets, thus the proportion of DirectSpec+’s running time to that of LightGCN/BPR becomes smaller. Meanwhile, SGL-ED is more computationally expensive on larger datasets, and DirectSpec+(LightGCN) is more efficient than it with more significant improvement over LightGCN. Furthermore, despite the additional training time per epoch added to the encoder, the scalability and utility have actually been improved. For instance, the required epochs are 250 and 600 for BPR and LightGCN on Gowalla, respectively, as opposed to just 80 and 60 epochs for DirectSpec(MF) and DirectSpec(LightGCN), resulting in 2.2x and 9.8x speed-up, respectively.
(a) AFDGCF on CiteULike
(b) LogDet on Gowalla
(c) DropEdge on Yelp
Figure 7. Comparison with several methods.
@@ 5.2.2. Efficiency
(c) Comparison of erank and average similarity on pairs in test sets
Figure 8. How DirectSpec prevents collapse (on Yelp).
5.2.3. Comparison with Negative Sampling
@@
The results in Table 3 demonstrates the superiority of DirectSpec over negative sampling based methods including BPR and CCL in terms of performance. Furthermore, we compare how these two types of methods alleviate collapse issue by optimizing them solely based on negative pairs. In Figure 6 (a) and (b), we observe that the erank of MF (BCE) consistently drops without optimization on the positive pairs, resulting in a complete collapse, and the erank drops more quickly as increasing the negative sampling ratios, whereas the erank of DirectSpec remains unchanged throughout the training. This observation indicates that there exists a dynamic balance between positive and negative samplings, the lack of either of them cannot prevent the embeddings from collapsing. Furthermore, since negative sampling is equivalent to a high pass filter, it can only alleviate the collapse caused by low pass filters (e.g., optimization on observed interactions), while DirectSpec is equivalent to an all pass filter that can tackle the collapse under any situations. Figure 6 (c) and (d) demonstrate that DirectSpec can recover the embeddings from collapsing caused by negative sampling.
5.2.4. Comparison with SSL-based methods
Particularly, we focus on AFDGCF (BT) and LogDet that can be considered as special cases of DirectSpec. We show how accuracy and erank change with the hyperparameters $\alpha$ and $\lambda$ controlling the intensity of spectrum balancing in Figure 7 (a) and (b), respectively. We observe that both performance and erank are not sensitive to $\alpha$ on AFDGCF, which are significantly inferior to DirectSpec+. While the erank consistently increases with $\lambda$ and reaches almost $d$ at $\lambda=0.9$ on LogDet, the accuracy drops early at $\lambda=0.3$ with erank=61.64. On the other hand, the best performance of DirectSpec+ is achieved at erank=63.77, showing significant improvement over LogDet. This observation indicates that DirectSpec+ can balance the spectrum in a more reasonable way, achieving a better trade-off between alignment and spectrum balancing than LogDet and AFDGCF.
5.2.5. Comparison with Solutions to Overs-Smoothing
[…]
(a) CiteULike
(b) Yelp
Figure 9. Impact of $\alpha$ on accuracy (nDCG@10) and erank.
(a) CiteULike
(b) Yelp
(c) Gwoalla
(d) Gowalla